# Learning to Refine Input Constrained Control Barrier Functions via Uncertainty-Aware Online Parameter Adaptation

### v1

@@ Abstract
Control Barrier Functions (CBFs) have become powerful tools for ensuring safety in nonlinear systems. However, finding valid CBFs that guarantee persistent safety and feasibility remains an open challenge, especially in systems with input constraints. Traditional approaches often rely on manually tuning the parameters of the class $\mathcal{K}$ functions of the CBF conditions a priori. The performance of CBF-based controllers is highly sensitive to these fixed parameters, potentially leading to overly conservative behavior or safety violations. To overcome these issues, this paper introduces a learning-based optimal control framework for online adaptation of Input Constrained CBF (ICCBF) parameters in discrete-time nonlinear systems. Our method employs a probabilistic ensemble neural network to predict the performance and risk metrics, as defined in this work, for candidate parameters, accounting for both epistemic and aleatoric uncertainties. We propose a two-step verification process using Jensen-Rényi Divergence and distributionally-robust Conditional Value at Risk to identify valid parameters. This enables dynamic refinement of ICCBF parameters based on current state and nearby environments, optimizing performance while ensuring safety within the verified parameter set. Experimental results demonstrate that our method outperforms both fixed-parameter and existing adaptive methods in robot navigation scenarios across safety and performance metrics.  [^1] Our project page can be found at: https://www.taekyung.me/online-adaptive-cbf
@@ I INTRODUCTION
Safety-critical optimal control is one of the fundamental challenges in robotics, e.g., robots need to navigate safely around obstacles while optimizing various criteria such as energy consumption. Control Barrier Functions (CBFs) have been developed in recent years as a tool to encode safety for nonlinear systems [1, 2, 3, 4, 5]. High Order CBFs extend this concept to systems with constraint functions of higher relative degree with respect to the system dynamics [6, 7, 8, 9]. However, a significant challenge lies in finding valid CBFs under input constraints [10]. Functions that do not provide persistent safety and feasibility guarantees are often referred to as candidate CBFs [11].
@@ II PRELIMINARIES
Given a valid discrete-time CBF $h$ satisfying condition (2) with the associated set $\mathcal{S}$, any control input ${\bm{u}}_{t}\in K_{\textup{cbf}}({\bm{x}}_{t})$, with $K_{\textup{cbf}}({\bm{x}}_{t})\coloneqq\{{\bm{u}}_{t}\in\mathcal{U}:\Delta h({\bm{x}}_{t},{\bm{u}}_{t})+\alpha(h({\bm{x}}_{t}))\geq 0\}$, renders the set $\mathcal{S}$ forward invariant for System (1). The proof can be found in [34, Theorem 1].
@@ III-A Input Constrained Control Barrier Functions
More importantly, the ICCBF structure explicitly addresses the problem of generating valid CBFs under input constraints, i.e., $\mathcal{U}\neq\mathbb{R}^{m}$. Concretely, suppose that the safe set $\mathcal{S}$ defined by $h$ cannot be rendered forward invariant by any feedback control input ${\bm{u}}\in K_{\textup{cbf}}({\bm{x}})$, since there exist some states where it requires ${\bm{u}}\notin\mathcal{U}$ to render safe.
@@
$\mathcal{C}_{1}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{0}({\bm{x}}_{t};\bm{\alpha})\geq 0\}=\mathcal{S}$ (4a)
$\mathcal{C}_{2}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{1}({\bm{x}}_{t};\bm{\alpha})\geq 0\}$ (4b)
@@
$\mathcal{C}_{r}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{r-1}({\bm{x}}_{t};\bm{\alpha})\geq 0\}.$ (4c)
@@
and it is conditioned on the class $\mathcal{K}$ functions $\bm{\alpha}$.
@@
Given the input constrained discrete-time system (1), if $b_{r}$ is a discrete-time ICCBF, then any control input ${\bm{u}}_{t}\in K_{\textup{iccbf}}({\bm{x}}_{t};\bm{\alpha})$, with $K_{\textup{iccbf}}({\bm{x}}_{t};\bm{\alpha})\coloneqq\{{\bm{u}}_{t}\in\mathcal{U}:\psi({\bm{x}}_{t},{\bm{u}}_{t};\bm{\alpha})\geq 0\}$, renders the set $\mathcal{C}^{*}(\bm{\alpha})\subseteq\mathcal{S}$ forward invariant. The proof is similar to [10, Theorem 1].
@@ III-B Problem Setup
satisfying local validity. Contrary, $\gamma_{0}=\gamma_{1}=0.5$ are not not locally valid since $\sup_{{\bm{u}}_{t}\in\mathcal{U}}\left[u_{t}\right]-1.75<0$. However, these class $\mathcal{K}$ functions can be locally valid at a different state, e.g., $x_{t}=-1,v_{t}=-1$, as the value becomes $\sup_{{\bm{u}}_{t}\in\mathcal{U}}\left[u_{t}\right]-0.75>0$.
@@ IV-B Training Data Generation - Offline Sampling
Based on the configuration above $(d_{0},\Delta\theta_{0},v_{0},\gamma_{0},\gamma_{1})$, we navigate the robot by solving the MPC-ICCBF problem (9) and record two metrics as the ground truth for prediction. To solve Problem 1, we compute the risk level $\phi\in\mathbb{R}$: $\phi=\max_{t\in[0,T_{\textup{max}}]}\Phi({\bm{x}}_{t},{\bm{u}}_{t})$, where $T_{\textup{max}}$ is the maximum simulation time. To approximately solve Problem 15, we compute the deadlock time $\delta\in\mathbb{R}$, which is an accumulated time segment where the robot exhibits a smaller velocity than a pre-defined threshold $v_{\textup{thr}}$. If the robot stops until $T_{\textup{max}}$, we set it as the maximum value $\delta=\delta_{\textup{max}}$. A training data pair consists of the input $X=[d_{0},\Delta\theta_{0},v_{0},\gamma_{0},\gamma_{1}]^{\top}$ and the ground truth $Y=[\phi,\delta]^{\top}$.
@@ IV-D2 Distributionally Robust Risk Assessment
Given that the predicted risk level $\hat{\phi}$ follows a GMM distribution due to ensemble predictions, we employ the distributionally robust Conditional Value at Risk (CVaR) as a risk metric. For notational simplicity, we slightly abuse the notation in this section by treating the risk level $\phi$ as the output vector $Y$. For $\hat{\phi}\sim{\bf E}_{b}(X)$, we define the Value at Risk (VaR) as:
@@ V-A Experimental Setup
For training data, we generate trajectories with $N=1$ obstacle in this paper. The robot and obstacle radii are set to $l_{\textup{robot}}=0.3$ m and $l_{\textup{obs}}=0.2$ m, respectively. The probability $\epsilon$ in CVaR constraint (29) is set to $0.05$. In total, we generate $M=16,808$ training samples offline. To simulate observation and model uncertainties, we add 3% of i.i.d. Gaussian noise to $d_{0},\Delta\theta_{0},v_{0}$ in the input data $X$. The PENN model ${\bf F}$ comprises $B=3$ ensemble models, each has 5 MLP layers with ReLU activation function. The rest of the implementation details follow [42]. More implementation details can be found in our public code repository. [^2] https://github.com/tkkim-robot/online_adaptive_cbf
@@ V-B Online Adaptive MPC-ICCBF
We compare the performance of our method with four different approaches: MPC-ICCBF using 1) “Low” ($\gamma_{0}=\gamma_{1}=0.01$) and 2) “High” ($\gamma_{0}=\gamma_{1}=0.2$) fixed class $\mathcal{K}$ function parameters, 3) Optimal-decay CBF-QP [28] that can optimize class $\mathcal{K}$ functions online (abbreviated as “Opt. QP”), and 4) “Optimal-decay MPC-CBF” [29] where it incorporates the same optimization technique as [28] within the MPC structure (abbreviated as “Opt. MPC”). For 3) and 4), we optimize both $\gamma_{0}$ and $\gamma_{1}$ and penalize them in the cost function. For our method, we empirically set the adaptation range as $\gamma_{0},\gamma_{1}\in(0.01,0.2)$.
@@
The conservative approach, using “Low” ICCBF parameters, is significantly affected by the initial nearby obstacle due to the ICCBF constraints. In some cases, the high initial velocities $v_{0}\in{\bm{x}}_{0}$ result in states that fall outside the inner safe set defined by fixed class $\mathcal{K}$ functions, leading to collisions. The aggressive approach with “High” ICCBF parameters becomes infeasible in early iterations and eventually collides with obstacles. While “Opt. QP” shows zero collisions, it is prone to deadlock. “Opt. MPC” avoids infeasibility but still results in collisions, as it only guarantees the point-wise feasibility of the solution rather than persistent safety. In contrast, our proposed method actively refines the class $\mathcal{K}$ function parameters online based on the PENN model’s predictions. It successfully reaches the goal without collisions or deadlocks, achieving the highest goal-reaching rate in both scenarios and demonstrating its ability to optimize performance while ensuring safety.

### v2

@@ Abstract
Control Barrier Functions (CBFs) have become powerful tools for ensuring safety in nonlinear systems. However, finding valid CBFs that guarantee persistent safety and feasibility remains an open challenge, especially in systems with input constraints. Traditional approaches often rely on manually tuning the parameters of the class $\mathcal{K}$ functions of the CBF conditions a priori. The performance of CBF-based controllers is highly sensitive to these fixed parameters, potentially leading to overly conservative behavior or safety violations. To overcome these issues, this paper introduces a learning-based optimal control framework for online adaptation of Input Constrained CBF (ICCBF) parameters in discrete-time nonlinear systems. Our method employs a probabilistic ensemble neural network to predict the performance and risk metrics, as defined in this work, for candidate parameters, accounting for both epistemic and aleatoric uncertainties. We propose a two-step verification process using Jensen-Rényi Divergence and distributionally-robust Conditional Value at Risk to identify valid parameters. This enables dynamic refinement of ICCBF parameters based on current state and nearby environments, optimizing performance while ensuring safety within the verified parameter set. Experimental results demonstrate that our method outperforms both fixed-parameter and existing adaptive methods in robot navigation scenarios across safety and performance metrics. [Project Page] [^1] Project page: https://www.taekyung.me/online-adaptive-cbf [Code] [Video]
@@ I INTRODUCTION
Safety-critical optimal control is one of the fundamental challenges in robotics, e.g., robots need to navigate safely around obstacles while optimizing various criteria such as energy consumption. Control Barrier Functions (CBFs) have been developed as a tool to encode safety for nonlinear systems and applied to high-level planning and multi-agent systems [1, 2, 3, 4, 5, 6, 7, 8, 9]. High Order CBFs extend this concept to systems with constraints of higher relative degree with respect to the system dynamics [10, 11, 12, 13]. However, a significant challenge lies in finding valid CBFs under input constraints [14]. Functions that do not provide persistent safety and feasibility guarantees are often referred to as candidate CBFs [15].
@@ II PRELIMINARIES
Given a discrete-time CBF $h$ satisfying condition (2) with the associated set $\mathcal{S}$, any control input ${\bm{u}}_{t}\in K_{\textup{cbf}}({\bm{x}}_{t})$, with $K_{\textup{cbf}}({\bm{x}}_{t})\coloneqq\{{\bm{u}}_{t}\in\mathcal{U}:\Delta h({\bm{x}}_{t},{\bm{u}}_{t})+\alpha(h({\bm{x}}_{t}))\geq 0\}$, renders the set $\mathcal{S}$ forward invariant for System (1). The proof can be found in [38, Theorem 1].
@@ III-A Input Constrained Control Barrier Functions
More importantly, the ICCBF structure explicitly addresses the problem of generating CBFs under input constraints, i.e., $\mathcal{U}\neq\mathbb{R}^{m}$. Concretely, suppose that the safe set $\mathcal{S}$ defined by $h$ cannot be rendered forward invariant by any feedback control input ${\bm{u}}\in K_{\textup{cbf}}({\bm{x}})$, since there exist some states where it requires ${\bm{u}}\notin\mathcal{U}$ to render safe.
@@
$\mathcal{C}_{0}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{0}({\bm{x}}_{t};\bm{\alpha})\geq 0\}=\mathcal{S}$ (4a)
$\mathcal{C}_{1}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{1}({\bm{x}}_{t};\bm{\alpha})\geq 0\}$ (4b)
@@
$\mathcal{C}_{r}(\bm{\alpha})$$\coloneqq\{{\bm{x}}_{t}\in\mathcal{X}:b_{r}({\bm{x}}_{t};\bm{\alpha})\geq 0\}.$ (4c)
@@
and it is dependent on the class $\mathcal{K}$ functions $\bm{\alpha}$.
@@
Given the input constrained discrete-time system (1), if $b_{r}$ is a discrete-time ICCBF, then any control input ${\bm{u}}_{t}\in K_{\textup{iccbf}}({\bm{x}}_{t};\bm{\alpha})$, with $K_{\textup{iccbf}}({\bm{x}}_{t};\bm{\alpha})\coloneqq\{{\bm{u}}_{t}\in\mathcal{U}:\psi({\bm{x}}_{t},{\bm{u}}_{t};\bm{\alpha})\geq 0\}$, renders the set $\mathcal{C}^{*}(\bm{\alpha})\subseteq\mathcal{S}$ forward invariant. The proof is similar to [14, Theorem 1].
@@ III-B Problem Setup
satisfying local validity. Contrary, $\gamma_{0}=\gamma_{1}=0.5$ are not locally valid since $\sup_{{\bm{u}}_{t}\in\mathcal{U}}\left[u_{t}\right]-1.75<0$. However, these class $\mathcal{K}$ functions can be locally valid at a different state, e.g., $x_{t}=-1,v_{t}=-1$, as the value becomes $\sup_{{\bm{u}}_{t}\in\mathcal{U}}\left[u_{t}\right]-0.75>0$.
@@ IV-B Training Data Generation - Offline Sampling
Based on the configuration above $(d_{0},\Delta\theta_{0},v_{0},\gamma_{0},\gamma_{1})$, we navigate the robot by solving the MPC-ICCBF problem (9) and record two metrics as the ground truth for prediction. To solve Problem 1, we compute the risk level $\phi\in\mathbb{R}$: $\phi=\max_{t\in[0,T_{\textup{max}}]}\Phi({\bm{x}}_{t},{\bm{u}}_{t})$, where $T_{\textup{max}}$ is the maximum simulation time. To approximately solve Problem 15, we compute the deadlock time $\delta\in\mathbb{R}$, which is an accumulated time segment where the robot exhibits a smaller velocity than a pre-defined threshold $v_{\textup{thr}}$. If the robot remains stationary until $T_{\textup{max}}$, we set it as the maximum value $\delta=\delta_{\textup{max}}$. A training data pair consists of the input $X=[d_{0},\Delta\theta_{0},v_{0},\gamma_{0},\gamma_{1}]^{\top}$ and the ground truth $Y=[\phi,\delta]^{\top}$.
@@ IV-D2 Distributionally Robust Risk Assessment
Given that the predicted risk level $\hat{\phi}$ follows a GMM distribution due to ensemble predictions and an analytical solution for its Conditional Value at Risk (CVaR) exists, we employ the distributionally robust CVaR as a risk metric. For notational simplicity, we slightly abuse the notation in this section by treating the risk level $\phi$ as the output vector $Y$. For $\hat{\phi}\sim{\bf E}_{b}(X)$, we define the Value at Risk (VaR) as:
@@ V-A Experimental Setup
For training data, we generate trajectories with $N=1$ obstacle in this paper. The robot and obstacle radii are set to $l_{\textup{robot}}=0.3$ m and $l_{\textup{obs}}=0.2$ m, respectively. The probability $\epsilon$ in CVaR constraint (29) is set to $0.05$. In total, we generate $M=16,808$ training samples offline. To simulate observation and model uncertainties, we add 3% of i.i.d. Gaussian noise to $d_{0},\Delta\theta_{0},v_{0}$ in the input data $X$. The PENN model ${\bf F}$ comprises $B=3$ ensemble models, each has 5 MLP layers with ReLU activation function. The rest of the implementation details follow [46]. More implementation details can be found in our public code repository.
@@ V-B Online Adaptive MPC-ICCBF
We compare the performance of our method with four different approaches: MPC-ICCBF using 1) “Low” ($\gamma_{0}=\gamma_{1}=0.01$) and 2) “High” ($\gamma_{0}=\gamma_{1}=0.2$) fixed class $\mathcal{K}$ function parameters, 3) Optimal-decay CBF-QP [32] that can optimize class $\mathcal{K}$ functions online (abbreviated as “Opt. QP”), and 4) “Optimal-decay MPC-CBF” [33] where it incorporates the same optimization technique as [32] within the MPC structure (abbreviated as “Opt. MPC”). For 3) and 4), we optimize both $\gamma_{0}$ and $\gamma_{1}$ and penalize them in the cost function. For our method, we empirically set the adaptation range as $\gamma_{0},\gamma_{1}\in(0.01,0.2)$. As demonstrated in [14], when the relative degree of the constraint function $h$ with respect to the control input matches the order to which the ICCBF is constructed, the ICCBF trivially reduces to an HOCBF in continuous time. Therefore, we adopt a discrete-time HOCBF implementation in our experiments.
@@
The conservative approach, using “Low” ICCBF parameters, is significantly affected by the initial nearby obstacle due to the ICCBF constraints. In some cases, the high initial velocities $v_{0}\in{\bm{x}}_{0}$ result in states that fall outside the inner safe set defined by fixed class $\mathcal{K}$ functions, leading to collisions. The aggressive approach with “High” ICCBF parameters becomes infeasible in early iterations and eventually collides with obstacles. While “Opt. QP” shows fewer collisions, it is prone to deadlock. “Opt. MPC” avoids infeasibility but still results in collisions, as it only guarantees the point-wise feasibility of the solution rather than persistent safety. In contrast, our proposed method actively refines the class $\mathcal{K}$ function parameters online based on the PENN model’s predictions. It successfully reaches the goal without collisions or deadlocks, achieving the highest goal-reaching rate in both scenarios and demonstrating its ability to optimize performance while ensuring safety.